# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio Choice

### v5

@@ 1 Introduction
In the remainder of this paper, we first present the multi-asset Merton problem with exogenous states in Section 2. Section 3 introduces our Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm, its Projected PG-DPO (P-PGDPO) extension, and provides a theoretical justification for the effectiveness of the P-PGDPO approach under specified regularity conditions. Numerical experiments presented in Section 4 confirm the framework’s scalability and near-optimal performance within this testbed, illustrating its potential for high-dimensional applications. Finally, Section 5 concludes and discusses potential future directions.
2 Multi-Asset Continuous-Time Portfolio Choice Problem with State Variables
@@
We consider the continuous-time portfolio choice problem introduced by Merton, (1973), in which a representative investor allocates wealth across $n$ risky assets and a risk-free asset, while consuming continuously over time. The investor selects an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}=(\pi_{1,t},\ldots,\pi_{n,t})^{\top}$, where $\pi_{i,t}$ denotes the proportion of current wealth $X_{t}$ invested in the $i$-th risky asset at time $t$ and notation denotes the transpose of a vector or a matrix.
The investment opportunities are characterized by an exogenous $k$-dimensional state process $\mathbf{Y}_{t}=(Y_{1,t},\ldots,Y_{k,t})^{\top}\in\mathbb{R}^{k}$, which governs the drift and volatility of asset returns. For instance, $\mathbf{Y}_{t}$ could include macroeconomic indices, volatility factors, exchange rates, etc. The investor’s optimization problem involves dynamically choosing both the portfolio weight vector $\bm{\pi}_{t}\in\mathbb{R}^{n}$ and the consumption rate $C_{t}$, conditional on the current state $\mathbf{Y}_{t}$ and wealth $X_{t}$.
To analyze this problem, we apply Pontryagin’s maximum principle (Pontryagin, , 2014; Yong and Zhou, , 2012), introducing costate variables corresponding to both the wealth process $X_{t}$ and the state process $\mathbf{Y}_{t}$. From a computational perspective, the control variables $\{\bm{\pi}_{t},C_{t}\}$ can be parameterized using neural networks, facilitating the implementation of policy gradient methods or DP algorithms in high-dimensional, state-dependent settings.
2.1 Problem Formulation with State Variables and Neural Network Controls
We assume that each component $Y_{i,t}$ of the state process $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ evolves according to an Itô diffusion:
@@ 2.1 Problem Formulation with State Variables and Neural Network Controls
where the Brownian motions $W^{Y}_{i,t}$ are potentially correlated, satisfying
$\operatorname{Cov}\bigl{(}dW^{Y}_{i,t},\,dW^{Y}_{j,t}\bigr{)}=\phi_{ij}\,dt,$
with constant correlations $\phi_{ij}$. In vector form, the state process satisfies
where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})\in\mathbb{R}^{k}$ is the drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\in\mathbb{R}^{k\times k}$ is a diagonal matrix of diffusion coefficients—i.e.,
We consider a financial market consisting of $n$ risky assets and one risk-free asset. The risk-free asset earns an instantaneous rate $r(t,\mathbf{Y}_{t})$, which may depend on time and the state variable. The price $S_{i,t}$ of the $i$-th risky asset evolves according to
@@
where $\mu_{i}(t,\mathbf{Y}_{t})$ is the expected return and $\sigma_{i}(t,\mathbf{Y}_{t})$ the volatility of asset $i$. The Brownian motions $W^{X}_{i,t}$ driving the risky asset dynamics satisfy
$\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt,\quad\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt,$
with constant correlations $\psi_{ij}$ and $\rho_{ij}$.
In vector notation, let $\bm{\mu}(t,\mathbf{Y}_{t})\in\mathbb{R}^{n}$ denote the vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})\in\mathbb{R}^{n\times n}$ the diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}\in\mathbb{R}^{n}$ the Brownian motion vector for the risky assets. Then the dynamics of the vector of risky asset prices $\mathbf{S}_{t}\in\mathbb{R}^{n}$ can be compactly written as
To clarify the structure of risk in the economy, we define the combined $(n+k)$-dimensional Brownian motion vector
$\mathbf{W}_{t}=\begin{pmatrix}\mathbf{W}^{X}_{t}\\ \mathbf{W}^{Y}_{t}\end{pmatrix}.$
The instantaneous covariance matrix of its increments $d\mathbf{W}_{t}$ is denoted by $\bm{\Omega}\,dt$, with
$\operatorname{Cov}(d\mathbf{W}_{t})=\bm{\Omega}\,dt=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}\,dt,$
where $\Psi=(\psi_{ij})\in\mathbb{R}^{n\times n}$ is the covariance matrix of the asset-specific shocks $d\mathbf{W}^{X}_{t}$, $\Phi=(\phi_{ij})\in\mathbb{R}^{k\times k}$ is the covariance matrix of the state shocks $d\mathbf{W}^{Y}_{t}$, and $\rho=(\rho_{ij})\in\mathbb{R}^{n\times k}$ captures the cross-covariation between the two sources of risk.
Throughout, we work on a complete probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ that supports a $(n+k)$-dimensional Brownian motion $\mathbf{W}_{t}$ and its natural filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$, augmented to satisfy the usual conditions (right-continuity and completeness).
The investor allocates wealth $X_{t}$ between the risk-free asset and the $n$ risky assets. The evolution of the investor’s wealth is described by the following stochastic differential equation (SDE):

### v6

@@ 1 Introduction
In the remainder of this paper, we begin with an introduction to the machine learning methods used for the PG-DPO framework in Section 2. Then we present the multi-asset Merton problem with exogenous states in Section 3. Section 4 introduces our Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm, its Projected PG-DPO (P-PGDPO) extension, and provides a theoretical justification for the effectiveness of the P-PGDPO approach under specified regularity conditions. Numerical experiments presented in Section 5 confirm the framework’s scalability and near-optimal performance within this testbed, illustrating its potential for high-dimensional applications. Finally, Section 6 concludes and discusses potential future directions of research.
2 Machine Learning Basics
Our proposed Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework is built upon fundamental concepts from machine learning. Before presenting our approach, this section introduces the computational methods and reinforcement learning principles that underpin our methodology. We begin with computational graphs and the algorithm of backpropagation through time (BPTT), which form the basis for efficient gradient calculation in sequential settings. We then discuss reinforcement learning and policy-based methods, highlighting their connection to dynamic portfolio choice and clarifying the distinctions between our framework and conventional RL approaches. Since our method assumes that the dynamics of the financial market are known, it can be classified as a form of model-based reinforcement learning.
2.1 Computational Graphs and Backpropagation Through Time
A computational graph provides a structured way to represent complex mathematical problems as a network of simple operations. In the portfolio optimization problem, the graph encodes the entire process, from the initial wealth and market states to the final objective function $J$, through sequential transformations of wealth and state variables. This representation is not merely a visualization but a computational tool: it enables us to systematically apply the chain rule of calculus across the entire problem.
Once we have this graph, we need an efficient way to understand how small changes in our decisions (the portfolio weights, parameterized by $\theta$) affect the final outcome. This is achieved through backpropagation, which works by starting from the final outcome and moving backward through the graph, systematically applying the chain rule at each step to calculate the contribution of each parameter to the final result. It is an algorithm that efficiently computes the gradient, or the sensitivity of the final output with respect to every parameter in the graph, denoted as $\nabla_{\theta}J$.
When applied to decision problems evolving over time, such as dynamic portfolio choice, the method is known as Backpropagation Through Time (BPTT). A key advantage of our model-based approach is that the dynamics of the environment, such as the stochastic differential equations (SDEs) for asset prices, are explicitly known. This allows us to compute exact analytical gradients for a given sample path through the simulation, rather than relying on estimation. BPTT is the core mechanism that enables our model to learn from simulated market trajectories by calculating explicit gradients. For a given trajectory, it computes the gradient of the final utility $J$ with respect to the policy parameters $\theta$ by propagating sensitivities backward through time. The influence of wealth in any intermediate time step, $X_{t+1}$, on the final result is found by applying the chain rule sequentially from the terminal state $X_{T}$:
$\frac{\partial J}{\partial X_{t+1}}=\frac{\partial J}{\partial X_{T}}\frac{\partial X_{T}}{\partial X_{T-1}}\frac{\partial X_{T-1}}{\partial X_{T-2}}\cdots\frac{\partial X_{t+2}}{\partial X_{t+1}}=\frac{\partial J}{\partial X_{T}}\prod_{k=t+1}^{T-1}\frac{\partial X_{k+1}}{\partial X_{k}}.$
This term represents the gradient ”flowing back” from the end of the simulation. The total gradient with respect to the policy parameters $\theta$ is then calculated by summing up the contributions from the policy $\pi_{t}$ at every time step, weighted by this backward propagated sensitivity:
$\frac{\partial J}{\partial\theta}=\sum_{t=0}^{T-1}\frac{\partial J}{\partial X_{t+1}}\frac{\partial X_{t+1}}{\partial\pi_{t}}\frac{\partial\pi_{t}}{\partial\theta}$
Since the dynamics of the financial market are explicitly modeled, our framework can compute analytical gradients along each sample path. This contrasts with model-free methods, which must rely on repeated trial-and-error estimates. In our method, BPTT thus plays a dual role: it provides the gradients necessary for optimizing the policy parameters, and at the same time, it generates pathwise costate estimates that are central to Pontryagin’s Maximum Principle, which will be introduced in Section 3.2.
2.2 Reinforcement Learning and Policy-Based Methods
Reinforcement Learning (RL) provides a general framework in which an agent learns to make decisions by interacting with an environment. At each point in time, the agent observes the current state, chooses an action, and receives a reward, with the goal of maximizing cumulative rewards over time. The mapping of this paradigm to dynamic portfolio choice is direct. The agent corresponds to an investor, the environment to a stochastic financial market, the state to the current wealth and economic factors, the action to portfolio allocation and consumption, and the reward to the utility derived from consumption and terminal wealth.
A central distinction in RL lies between value-based and policy-based methods. Value-based approaches estimate value functions and derive actions indirectly, whereas policy-based methods directly optimize a parameterized policy $\pi_{\theta}$. In our setting, the policy is represented by a neural network with parameters $\theta$, which maps the current state $s_{t}$ (e.g., market conditions, current wealth) to an action $a_{t}=\pi_{\theta}(s_{t})$ (e.g., a specific portfolio allocation). The optimization problem reduces to finding the parameter set $\theta^{*}$ that maximize the expected total reward (our objective function $J$):
$\theta^{*}=\arg\max_{\theta}\mathbb{E}[J(\theta)].$
Parameter updates follow the stochastic gradient ascent rule,
$\theta_{k+1}=\theta_{k}+\alpha\mathbb{E}\left[\frac{\partial J(\theta_{k})}{\partial\theta}\right]$
where $\alpha$ is a learning rate and gradients are computed through BPTT.
While our PG-DPO framework is conceptually rooted in this class of algorithms, it departs in important ways. First, it is a model-based RL method. Unlike model-free approaches such as Proximal Policy Optimization or actor-critic schemes, which require repeated trial-and-error exploration to estimate gradients, as seen in Dai et al., (2025) that tackle the Merton problem, our framework exploits the explicitly known dynamics of financial markets. This allows us to compute exact analytical gradients through BPTT, leading to both greater efficiency and higher numerical accuracy.
Second, our approach is explicitly guided by Pontryagin’s Maximum Principle (PMP). Standard RL methods often converge to policies that deliver nearly optimal utility values but may diverge substantially from the structural form of the true solution. In other domains, such as Atari, Go, or autonomous driving, such structural fidelity is less critical, since achieving a near-optimal utility value alone is often sufficient for strong performance. In dynamic portfolio choice, however, the decomposition of the optimal strategy into myopic demand and intertemporal hedging demand is both theoretically central and practically important. By leveraging PMP, our framework ensures that both components are faithfully captured, yielding policies that are structurally consistent with financial theory.
A further distinction arises in comparison with Duarte et al., (2024), whose method employs an actor–critic scheme in which the critic is trained by minimizing the HJB residual in a PINN-like fashion. [^4] The HJB residual is commonly defined to be the cumulative squared deviation of the left-hand side of the HJB equation from 0.. However, this approach is designed exclusively for infinite-horizon problems, where the optimal solution is stationary and time-independent. Although appropriate for certain theoretical settings, this restriction makes it less suitable for most practical applications, such as retirement planning, endowment management, or finite-maturity products. By contrast, our PG-DPO framework is explicitly tailored to the finite-horizon setting, incorporating time-to-maturity as a state variable. This enables the algorithm to capture horizon effects, including the time-varying structure of hedging demand and the transition of optimal strategies as maturity approaches. Moreover, since the long-horizon case can be treated as a special instance of finite-horizon formulation, our framework naturally extends to settings such as those analyzed in Section 5.3, where the horizon is sufficiently long to yield stationary policies. This flexibility is in sharp contrast to the approach of Duarte et al., (2024), which is limited to the infinite horizon case and cannot easily accommodate horizon-dependent dynamics that are crucial in practice.
In sum, our method combines the principles of reinforcement learning with the structural discipline of PMP, while extending applicability to finite-horizon continuous-time finance. This integration ensures that learned policies are not only utility-maximizing but also theoretically well-structured, marking a significant departure from existing RL approaches.
3 The Model and Pontryagin’s Maximum Principle
3.1 General Multi-Asset Model
@@
We consider the continuous-time portfolio choice problem introduced by Merton, (1973), in which a representative investor allocates wealth between risky assets and a risk-free asset, while continuously consuming over time. [^5] While we focus on Brownian motion as the source of return risk, the methods can be extended to accommodate jumps. Furthermore, if a discrete-time model is viewed as an approximation of a continuous-time model, our approach can be adapted accordingly.
The stochastic investment opportunities are characterized by an exogenous $k$-dimensional state process $\mathbf{Y}_{t}=(Y_{1,t},\ldots,Y_{k,t})^{\top}\in\mathbb{R}^{k}$, where notation denotes the transpose of a vector (or a matrix), and each component $Y_{i,t}$ evolves according to an Itô diffusion:
@@
where the Brownian motions $W^{Y}_{i,t}$ are potentially correlated, satisfying $\operatorname{Cov}\bigl{(}dW^{Y}_{i,t},\,dW^{Y}_{j,t}\bigr{)}=\phi_{ij}\,dt,$ with constant correlations $\phi_{ij}$. For instance, $\mathbf{Y}_{t}$ could include macroeconomic indices, volatility factors, exchange rates, etc.
We consider a financial market consisting of $n$ risky assets and one risk-free asset. The risk-free asset earns an instantaneous rate $r(t,\mathbf{Y}_{t})$, which depends on time and the state variable. The price of the $i$-th risky asset, $S_{i,t}$, evolves according to
@@
where $\mu_{i}(t,\mathbf{Y}_{t})$ is the expected return and $\sigma_{i}(t,\mathbf{Y}_{t})$ the volatility. Brownian motions $W^{X}_{i,t}$ have correlations that satisfy $\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt$, and $\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt,$ with constant correlations $\psi_{ij}$ and $\rho_{ij}$. The investor selects an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}=(\pi_{1,t},\ldots,\pi_{n,t})^{\top}$, where $\pi_{i,t}$ denotes the proportion of current wealth $X_{t}$ invested in the $i$-th risky asset at time $t$.
To clarify the structure of risk in the economy, we define the combined $(n+k)$-dimensional Brownian motion vector $\mathbf{W}_{t}=\begin{pmatrix}\mathbf{W}^{X}_{t}\\ \mathbf{W}^{Y}_{t}\end{pmatrix}$. where $\mathbf{W}_{t}^{X}$ and $\mathbf{W}_{t}^{Y}$ are Brownian vectors for the risky assets and state variables. Then, the instantaneous covariance matrix of its increments $d\mathbf{W}_{t}$ is denoted by $\bm{\Omega}\,dt$, with $\operatorname{Cov}(d\mathbf{W}_{t})=\bm{\Omega}\,dt=\begin{pmatrix}\Psi&\bm{\rho}\\ \bm{\rho}^{\top}&\Phi\end{pmatrix}\,dt,$ where $\Psi=(\psi_{ij})\in\mathbb{R}^{n\times n}$ is the correlation matrix of the asset-specific shocks $d\mathbf{W}^{X}_{t}$, $\Phi=(\phi_{ij})\in\mathbb{R}^{k\times k}$ is the correlation matrix of the state shocks $d\mathbf{W}^{Y}_{t}$, and $\bm{\rho}=(\rho_{ij})\in\mathbb{R}^{n\times k}$ captures the cross-correlation between the two sources of risk.
Throughout the paper, we consider a complete probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ that supports a standard $(n+k)$-dimensional Brownian motion $\mathbf{W}_{t}$ and its natural filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$, augmented to satisfy the usual conditions such as right-continuity and completeness. Then the investor’s wealth dynamics is given by